Question 8. Results of a multiple linear regression on these four variables are depicted in the table below.

Multiple Linear Regression: Summary Output

Coefficients

Intercept

-35.2

Training

2.5

Performance Score

8.2

Years of Service

1.1

a) Use this information to develop and interpret a Regression Equation.

b) Estimate the value of the dependent variable when an employee has attended 5 hours of training, receives a performance evaluation score of 3.6, and has been working for the company for 9 years.
Answer #1

a)

Regression equation

y = -35.2 + x1*Rtraininghours + x2*RPerfromance score +x3*RYears of service

y = -35.2 + x1*2.5 + x2*8.2 +x3*1.1

Interpretation

Multiple regression indicates effect of multiple factors on dependent variable

Rtraining, RPerfromance score,RYears of service are partial regression coefficients
x1,x2,x3 are independent variables

In multiple regression, regression coefficients are interpreted as change in dependent variable with 1 unit change of the independent variable considering all other variables constant

So, Change in dependent variable is 2.5 when training hours is changed by 1
Change in dependent variable is 8.2 when Performance score is changed by 1
Change in dependent variable is 1.1 when Years of service is changed by 1

b)

Value of dependent variable = -35.2 + 2.5*5 + 8.2*3.6 +1.1*9 = 16.72